OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1). A.) y = 1/3x − 3 B.) y = 1/3x + 17 C.) y = −3x − 3 D.) y = −3x + 17

OpenStudy (anonymous):

Can you find the slope of the line that you are given?

OpenStudy (anonymous):

I found it by putting 3x+y=7 into y=mx+b form.

OpenStudy (haseeb96):

y=7-3x here , b= 7 and m= -3 but we know that in perpendicular position we use m1m2=-1 m -3 =-1 m=1\3 so m=1\3 and the line passes from the (6,-1) (x,y)=(6,-1) y=mx+b -1=6(1\3)+b -1= 2+b b=-1-2 b=-3 now put it in the slope intercept from y=1\3 x -3 this is the true answer

OpenStudy (haseeb96):

A is the correct answer A) y=1\3 x -3